6. For a holiday party, Marcus is making a seven-layer torte, which consists of seven 9-inch diameter cake layers with a cream topping on each layer. He knows that one recipe of cream will cover 65 square inches of cake. How many batches of the topping must he make to complete his torte? (Use p = 3.14 as an approximate value)
thank you most graciously for your answers.
AD

Answer by checkley77(12844)

area=pir^2
3.14*4.5^2=63.585 in^2 per layer.
7*63.585=445.095 in^2 for all 7 layers.
445.095/65=6.85 batches of topping will be needed.